A solid rigid body rotating about an axis has a mass of 20kg and a radius of 2m from the axis of rotation.if its angular velocity is 25 rad/s,calculate its rotational kinetic energy

1
Expert's answer
2022-01-21T08:25:59-0500

"K=\frac 12 mv^2=\frac 12m\omega^2r^2=25~kJ."